OpenStudy (itrymath):

Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (itrymath):

@Directrix

OpenStudy (itrymath):

i thought of an equation and some points but not sure if it satisfies my equation

Directrix (directrix):

Do this and then I can help: Create your own equation written in standard form.

OpenStudy (itrymath):

y=5x-9

OpenStudy (itrymath):

i got some points (-1,4)

Directrix (directrix):

The Standard Form for a linear equation in two variables, x and y, is usually given as Ax + By = C where, if at all possible, A, B, and C are integers, and A is non-negative, and, A, B, and C have no common factors other than 1. ----------------------------------- Look in your text and see how standard form of a line is presented in your text.

Directrix (directrix):

y=5x-9 Slope/intercept form. It may be considered standard in your book. We want to get off to a good start. So, check to see how your text defines standard form of a line.

OpenStudy (itrymath):

actually i got that equation from another open study -_-

OpenStudy (itrymath):

liol

OpenStudy (itrymath):

https://prezi.com/2kuanyteuuw_/304-module-3-quiz/

OpenStudy (itrymath):

its helped on some problems but on this one it's to straight forward

Directrix (directrix):

Does your textbook define standard form of the equation of a line?

OpenStudy (itrymath):

no it doesn't

Directrix (directrix):

Working the problem won't be difficult. If only I could get you to tell me how your book defines standard form of a linear equation.

OpenStudy (itrymath):

it doesn't so just tell me how you define it

OpenStudy (itrymath):

my teacher talked about Ax+Bx=C

Directrix (directrix):

The standard form of an equation is Ax + By = C. In this kind of equation, x and y are variables and A, B, and C are integers. Okay, let's go with what the teacher said. First, we will need to write y = 5x -9 in standard form.

OpenStudy (itrymath):

okay how do we do it

Directrix (directrix):

y = 5x - 9 -y -y ---------- 0 = 5x - y - 9 +9 +9 ------------- 9 = 5x - y 5x -y = 9 ---> standard form (do you agree)

OpenStudy (itrymath):

yes very

Directrix (directrix):

If you agree, then this solution begins with the equation: 5x -y = 9 What is the first step now that we have the equation in standard form?

OpenStudy (itrymath):

+y

Directrix (directrix):

Determine any point that is a solution and will be on this line. Justify your answer algebraically. Where are the coordinates of the point you posted are on this line?

OpenStudy (itrymath):

wait no!

OpenStudy (itrymath):

-5x from both sides

OpenStudy (itrymath):

how do i get the right coordinates or they can be random?

Directrix (directrix):

You posted this point earlier in the thread: >> i got some points (-1,4)

OpenStudy (itrymath):

yes but that was random

Directrix (directrix):

Let's see if the point is on the line. Is (-1,4) on the line? 5x -y = 9 x = -1 and y = 4 Does 5* (-1) - 4 = 9 ? yes or no

OpenStudy (itrymath):

it equals -9=9 so no?

Directrix (directrix):

5* (-1) - 4 = -5 -4 = -9 which is not on the line We need another point.

Directrix (directrix):

In the equation if x = 0, what is the value of y that goes with that?

Directrix (directrix):

5x -y = 9 5*0 - y = 9 ---> Solve for y

OpenStudy (itrymath):

how about 2,2?

Directrix (directrix):

(2,2) is not on the line. Please do this: 5*0 - y = 9 ---> Solve for y

OpenStudy (itrymath):

-y=9

OpenStudy (itrymath):

so -9?

Directrix (directrix):

Yes. ( 0. -9) is a point on the graph of 5x -y = 9 because 5*0 - (-9) = 9

Directrix (directrix):

Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. The x-intercept is on the x-axis. All points on the x-axis have a y-coordinate of 0. To find the x-intercept, let y = 0 in the equation 5x -y = 9. 5*x - 0 = 9 5x = 9 x = 9/5 = 1.8 ( 1.8 , 0 ) for the x-intercept.

OpenStudy (itrymath):

okay but how come your way is different from hers? https://prezi.com/2kuanyteuuw_/304-module-3-quiz/

Directrix (directrix):

The video will not load. If you want to work with "her," then I will leave you to do that.

OpenStudy (itrymath):

@563blackghost

OpenStudy (leenathan):

this is alot i cant comprehend and also its the week end no ones on here much

OpenStudy (itrymath):

@sammixboo

OpenStudy (itrymath):

@Jamierox4ev3r

OpenStudy (leenathan):

@Austin.L

OpenStudy (itrymath):

you know how to do this?

OpenStudy (itrymath):

@mathmate

OpenStudy (itrymath):

@sweetburger

OpenStudy (itrymath):

@CaNdI5a18

OpenStudy (itrymath):

hey austin

OpenStudy (itrymath):

nevermind